OpenStudy (anonymous):

whats the correct simplfication 9y^0?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

can u answer these other questions for me?

OpenStudy (anonymous):

9 * y^0 = 9 * 1 = 9

jhonyy9 (jhonyy9):

9*1=9